There where 196 students that participated and I clean up event at a nearby lake is one fourth of the entire student population right and solve an equation to find the number of students at the school
Answer by Theo(13342)

if 196 students equal 1/4 of the student population, then the number of the students in the school is 196 / (1/4) = 196 * (4/1) = 196 * 4 = 784.

you would solve this algebraically by doing the following.

let x = number of students in the school.

since 1/4 of the student in the school is equal to 196, you get 1/4 * x = 196

multiply both sides of this equation by 4 to get x = 196 * 4

simplify by performing the multiplication to get x = 784